# Jaike van Twiller1    Yossiri Adulyasak2    Erick Delage2    Djordje Grbic1    Rune Møller Jensen1

### v3

@@ 6.3 Managerial Insights
Figure 4 performs a sensitivity analysis to examine the effects of SMIP scenario size and demand uncertainty, reporting with an average and 95% confidence interval (CI).
Value of Information. Figure 4(a) compares performance under non-anticipation, imperfect, and perfect information. The predictive accuracy of SMIP models improves with the number of scenarios, as shown by the stabilizing yet decreasing objective values. With 28 scenarios, perfect information enhances profit by 60-65%, while imperfect information via DRL with VP achieves a 25-40% improvement, emphasizing the role of information availability for decision quality.
Computational Time. While inference time is critical, training time remains relevant. Figure 4(b) illustrates the total computational time of 30 instances, with training time included for DRL. Computational cost increases significantly with the number of scenarios, with SMIP-NA exhibiting exponential growth and SMIP-PI showing steady growth. Solving the most accurate SMIP-NA takes approximately 2 hours per instance, which is intractable for stowage planners. In contrast, the DRL approach requires offline training of about 11 hours, after which solutions can be constructed in seconds.
@@
Adaptiveness. Figure 4(c) illustrates how policies adapt to variations in the spread of the demand distribution of unseen instances. As variability and uncertainty increase, both SAC and PPO policies achieve higher profits, demonstrating their ability to leverage uncertainty. Despite fluctuations in variability, both policies consistently generate feasible solutions.
Implications for Practical Deployment. Our findings are based on simulations, limited by the lack of real-world data. To ensure effective deployment, practitioners should focus on collecting and generating representative cargo demand data.
4444888812121212161616162020202024242424282828281,00010001{,}0001 , 0001,20012001{,}2001 , 2001,40014001{,}4001 , 4001,60016001{,}6001 , 6001,80018001{,}8001 , 800Number of ScenariosProfit (\$)SMIP-NASMIP-PI PPO/VPSAC/VP
(a) Profit with 95% CI across scenario sizes on 30 instances.
@@
(b) Total computational time across scenario sizes of 30 instances.
4444888812121212161616162020202024242424282828281,00010001{,}0001 , 0001,20012001{,}2001 , 2001,40014001{,}4001 , 4001,60016001{,}6001 , 6001,80018001{,}8001 , 8000.10.10.10.10.30.30.30.30.50.50.50.50.70.70.70.70.90.90.90.91,30013001{,}3001 , 3001,40014001{,}4001 , 4001,50015001{,}5001 , 500Coefficient of Variation (CV)Profit (\$)PPO/VPSAC/VPSMIP-NASMIP-PI PPO/VPSAC/VP
(c) Profit with 95% CI across CV levels on 30 unseen instances.
Figure 4: Sensitivity analysis of scenario size and demand spread.
7 Conclusion and Future Directions
This work introduces a novel MDP formulation for the MPP under demand uncertainty, incorporating realistic inequality constraints. We train an AM policy using actor-critic DRL methods with differentiable feasibility projections to construct MPP solutions. Experimental results demonstrate that our policy efficiently generates adaptive and feasible solutions, significantly outperforming baseline DRL methods and the SMIP-NA. This approach establishes an AI-driven decision-support policy for planning under uncertainty in a critical part of the global supply chain. Future work will extend the MPP formulation and scale to larger vessels and longer voyages, further enhancing the representativeness.
@@ 7 Conclusion and Future Directions
Human oversight is critical in decision-support tools to ensure fairness, accountability, and transparency. While models can enhance decision-making, they must not replace human judgment, particularly in high-stakes applications. We strongly encourage safeguards to mitigate automation bias, ensuring that users can override or challenge outputs, with continuous validation and auditing to uphold trust and fairness.
@@ Appendix A MDP of Master Planning Problem
A.1 Sets and Parameters
Provided the sets and parameters in Section 2, we introduce additional subsets of the transport set TR given port $p\in{P}$:
@@ A.1 Sets and Parameters
  • Arrival transports: $\textit{TR}^{\textit{AC}}_{p}=\{(i,j)\!\in\!{P}^{2}\!\mid\!i<p,j>p\}$
@@
For each combination $(i,j,k)$, we associate the expected demand $\mu^{(i,j,k)}$, standard deviation $\sigma^{(i,j,k)}$, TEU per container $\textit{teu}^{(i,j,k)}$, container weight $w^{(i,j,k)}$, and revenue per container $\textit{rev}^{(i,j,k)}$.
@@
The parameters $\mu$ and $\sigma$ are randomly generated, as shown in Appendix C.
A.2 Formal MDP
We define the MDP by decomposing the traditional CO problem outlined in Van Twiller et al. [2024a].
@@ A.3 Decomposed MDP
Algorithm 2 Episode of Augmented MDP
0:  $\mathcal{T}$, $\pi_{\theta}$, $\zeta$, $\mathcal{Q}$
1:  $q_{T_{\textit{seq}}}^{(i,j,k)}\sim\mathcal{Q}(\mu^{(i,j,k)},\sigma^{(i,j,k)})\;\forall(i,j)\in\textit{TR},k\in K$
@@
$T_{\text{leave port}}$ =$=\Big{\{}t\in H\ \mid\exists p\in P^{N_{P}-1}_{1}\text{ such that }$
@@
$T_{\text{new port}}$ =$=\Big{\{}t\in H\ \mid\exists p\in P^{N_{P}-1}_{1}\text{ such that }$
@@ A.3.2 Feasible Region
A.3.3 Substituting Load Operations for Actions
@@
A.3.4 Feasible Region for Actions
@@ A.4.2 Capacity Constraints
$^{(\textit{pol}_{t},\textit{pod}_{t},k_{t})}\mathbf{1}^{\top}x_{t}\leq c-u^{\prime}_{t}\textit{teu}\;,$ teu ,
A.4.3 Stability Constraints
@@
$\mathbf{1}^{\top}(\textit{lm}\odot u^{+}_{p})+\mathbf{1}^{\top}(\textit{lm}\odot u^{\prime}_{p})$ $\leq\overline{\textit{lcg}}\mathbf{1}^{\top}(w\odot u^{+}_{p})$
@@ A.5 Auxiliary Variables
$\textit{ho}(s,p)$ $=\textit{hm}(s,p)\Bigg{(}\sum_{k\in K}\sum_{\textit{tr}\in\textit{TR}^{\textit{AC}}_{p}}u_{t}^{(b,d^{\textit{above}},k,\textit{tr})}\Bigg{)}$ (15)
@@
${\textit{cm}}(s,p)$ $=\max\Bigg{(}\overline{\textit{cm}}(s,p)-\sum_{d\in D}\sum_{k\in K}\sum_{\textit{tr}\in\textit{TR}^{M}_{p}}u_{t}^{(0:|B|-1,d,k,\textit{tr})}$
$\qquad\qquad+u_{t}^{(1:|B|,d,k,\textit{tr})},0\Bigg{)}$ (17)
@@ Appendix B Feasibility Mechanisms
Table 2 provides an overview of implemented feasibility mechanisms.
Type Implementation Constraints
FR Composite loss Constraints $\textit{PH}(s_{t})$
VP $\textit{VP}(x_{t},A(s_{t}),b(s_{t}),\alpha_{v},\delta_{v})$ Constraints $\textit{PH}(s_{t})$
WS $\mathcal{W}(x_{t},q_{t})$ Demand $q_{t}$
PC $\mathcal{C}(x_{t},0,c-u^{\prime}_{t}\textit{teu})$ TEU capacity $c$
Table 2: Feasibility mechanisms and relation to constraints
B.1 Log Probability Adjustments
This subsection provides technical details on the adjustments made to the log-probability distribution of the policy as a result of non-linear transformations to distribution samples.
Projecting actions alters the policy’s probability density, necessitating consideration of the change of variables principle Bishop [2006]. This principle ensures valid volume scaling by requiring the transformation $f(x)$ to satisfy:
@@ B.1 Log Probability Adjustments
B.1.1 Weighted Scaling Projection Layer.
@@
B.1.2 Violation Projection Layer.
@@
Finally, we can confirm that Jacobian adjustment is allowed for the violation projection layer by:
@@
B.1.3 Policy Clipping
We can implement a clipped Gaussian distribution that enforces element-wise bounds on a standard Gaussian Fujita and Maeda [2018]. Let $\mu_{\theta}$ and $\sigma^{2}_{\theta}$ denote the policy’s mean and variance, with bounds $\textit{lb}_{\textit{pc}}$ and $\textit{ub}_{\textit{pc}}$, and $\Phi(\cdot)$ being the cumulative distribution function of the standard Gaussian. Actions are sampled from $\mathcal{N}(\mu_{\theta},\sigma^{2}_{\theta})$ and clipping the result to $[\textit{lb}_{\textit{pc}},\textit{ub}_{\textit{pc}}]$. Provided this transformation, we compute the log probabilities $\log\pi(x|s)$ for action $x$ by:
@@ B.1.3 Policy Clipping
B.2 Violation Projection Layer
@@
  • $\eta_{v}\in(0,1)$ is a step size parameter.
@@
Since $\mathcal{V}(x)$ is an elementwise projection onto nonnegative values, and some function $h(y)=\max(0,y)$ is convex and non-decreasing, then the function $g(x)=\|\mathcal{V}(x)\|_{2}^{2}$ is convex when $Ax-b$ is affine.
@@
By the standard descent lemma Boyd and Vandenberghe [2004], for a sufficiently small step size $\eta_{v}$, we have:
@@
Since $g(x_{k})$ is also lower-bounded by 00, it must converge to some limit $g^{*}\geq 0$. This implies:
@@ Appendix C Instance Generator
To ensure similar mean and variance of the instances, we derive parameters $a$ and $b$ from the definition of the continuous uniform distribution, as follows:
@@ Appendix D Multi-stage Stochastic MIP
D.1 Multi-stage Scenario Tree
@@
    A probability measure $P:V_{\textit{ST}}\to[0,1]$ assigning probabilities to nodes, ensuring:
@@
D.2 MIP Formulation
@@
In addition, we introduce the big $M$ as a sufficiently large constant used to enforce logical constraints.
Objective. The objective function (18) maximizes the revenue and minimizes hatch-overstowage and crane moves costs over scenario paths $\phi\in\mathcal{Z}$ each with probability $\mathbb{P}_{\phi}$. We assume each scenario path has equal probability.
@@
max $\sum_{\phi\in\mathcal{Z}}\mathbb{P}_{\phi}\sum_{p\in{P}}\sum_{b\in{B}}\sum_{d\in{D}}\sum_{k\in{K}}\sum_{\textit{tr}\in\textit{TR}^{\textit{+}}(p)}f_{1}\tilde{u}^{b,d,\phi}_{\textit{tr},k}$
$\qquad\qquad\qquad\qquad\qquad-f_{2}\tilde{\textit{ho}}^{\phi}_{p,b}-f_{3}\tilde{\textit{cm}}^{\phi}_{p}$ (18)
@@
$\sum_{k\in{K}}\sum_{\textit{tr}\in\textit{TR}^{\textit{AC}}(p)}\tilde{u}^{b,d_{o},\phi}_{\textit{tr},k}-M(1-\tilde{\textit{hm}}^{\phi}_{p,b})\leq\tilde{\textit{ho}}^{\phi}_{p,b}$
@@
${\overline{z}}^{\phi}_{p}-\sum_{b\in{b^{\prime}}}\sum_{d\in{D}}\sum_{k\in{K}}\sum_{\textit{tr}\in\textit{TR}^{M}(p)}\tilde{u}^{b,d,\phi}_{\textit{tr},k}\leq\tilde{\textit{cm}}^{\phi}_{p}$
@@ Appendix E Deep RL Implementation Details
To prevent confusion, we
E.1 PPO Algorithm
PPO is an on-policy reinforcement learning algorithm that seeks to maximize expected cumulative reward while enforcing stable policy updates via clipped importance sampling Schulman et al. [2017], as outlined in Algorithm 3. The agent collects trajectories, computing $n_{\text{ppo}}$-step return to evaluate performance with $V_{\theta}(s)$ as estimated state value:
@@ E.1 PPO Algorithm
To reduce variance, we adopt Generalized Advantage Estimation (GAE) Schulman et al. [2016]:
@@
E.2 SAC Algorithm
Soft Actor-Critic (SAC) is an off-policy reinforcement learning algorithm that optimizes both reward maximization and entropy to encourage efficient exploration Haarnoja et al. [2018], as outlined in Algorithm 4. It is based on maximum entropy reinforcement learning, which aims to learn a stochastic policy that not only maximizes cumulative rewards but also maintains high entropy for robustness and stability. SAC leverages a soft Q-learning approach, using two Q-functions to mitigate overestimation bias, an entropy-regularized policy update, and an automatically adjusted temperature parameter to balance exploration and exploitation.
@@ E.2 SAC Algorithm
1:  for each step $t$ do
2:     Sample action $x_{t}\sim\pi_{\theta}(x_{t}|s_{t})$
3:     Perform transition $s_{t+1}\sim\mathcal{T}(s_{t+1}|s_{t},x_{t})$
4:     Observe reward $r_{t}=\mathcal{R}(s_{t},x_{t})$,
5:     Store $(s_{t},x_{t},r_{t},s_{t+1})$ in $\mathcal{D}$.
6:  end for
7:  for each gradient update do
8:     Sample a minibatch $(s_{t},x_{t},r_{t},s_{t+1})$ from $\mathcal{D}$.
9:     Compute target Q-value: $Q_{\text{target}}(s_{t},x_{t})$
10:     Update parameters:  $\theta_{\text{critic}}^{l}\leftarrow\theta_{\text{critic}}^{l}-\eta_{c}\nabla_{l}\mathcal{L}_{\text{critic}}(\theta)\text{ for }l\in\{1,2\}$$\theta_{\text{actor}}\leftarrow\theta_{\text{actor}}-\eta_{a}\nabla\mathcal{L}_{\text{actor}}(\theta)$$\alpha\leftarrow\alpha-\eta_{\alpha}\nabla\mathcal{L}_{\alpha}(\theta)$$\theta_{\text{target}}^{l}\leftarrow\tau\theta_{\text{critic}}^{l}+(1-\tau)\theta_{\text{target}}^{l}\text{ for }l\in\{1,2\}$
11:  end for
E.3 Hyperparameters
Parameters of the MPP environment are shown in Table 3.
@@ E.3 Hyperparameters
Table 3: Environment parameters.
Table 4 provides the hyperparameters of projected and vanilla PPO and SAC.
Settings Projection Algorithms Vanilla Algorithms
Hyperparameters Symbol PPO SAC PPO SAC
@@
Number of Heads 8 8 4 4
Hidden Layer Size 128 128 256 256
Encoder Layers 3 3 2 1
Decoder Layers 3 3 3 3
@@
Dropout Rate $0.009$ $0.009$ $0.073$ $0.164$
Max Policy Std. $1.931$ $1.931$ $1.118$ $1.779$
@@
Learning Rate $\eta$ $2.04\times 10^{-4}$ $2.04\times 10^{-4}$ $9.64\times 10^{-4}$ $9.10\times 10^{-4}$
Batch Size 64 64 64 64
Embedding Size 128 128 128 128
@@
Feasibility Penalty $\lambda_{f}$ $0.0677$ $0.0677$ $0.302$ $0.065$
@@
Mini-batch Size 16 16 32 16
Update Epochs 5 1 1 1
Entropy Target N/A $-|{X}|$ N/A $-|{X}|$
Projection Learning Rate $\eta_{v}$ $0.05$ $0.05$ N/A N/A
Projection Epochs $100$ $100$ N/A N/A
Inference Projection Stop $\delta_{v}$ $100$ $100$ N/A N/A

### v5

@@ Introduction
Figure 5 performs sensitivity analyses on SMIP scenario size, demand uncertainty, and voyage length, reporting averages and 95% confidence intervals with $N=30$. We use the same hyperparameters as Table 2 unless stated otherwise.
Value of Information.
Figure 5(a) compares profits under non-anticipation, imperfect, and perfect information. As scenario sizes grow, the SMIP predictive accuracy improves and profits stabilize. At 28 scenarios, perfect information increases profits by 60–65% over non-anticipation, while DRL with imperfect information improves profits by 30–40%, showing the impact of information on decision quality.
Computational Time.
While inference time is crucial, training time also matters. Figure 5(b) shows computational time for 30 instances. SMIP-NA’s time grows exponentially with scenarios, reaching 2 hours per instance at 28 scenarios, while SMIP-PI grows steadily. SAC trains offline in 1–3 hours, and PPO takes 1–12 hours, except for AM-P with CP, which requires significantly longer. Inference is near-instant for all. Amortized, SAC and PPO average 8 and 10 minutes per instance, making DRL more efficient than SMIP-NA.
Adaptiveness.
@@
Figure 5(c) shows how policies adapt to demand uncertainty by varying the coefficient of variation across unseen instances. As variability and uncertainty increase, both SAC and PPO policies achieve higher average profits by exploiting demand outliers, demonstrating their ability to leverage uncertainty. Despite fluctuations in variability, both policies consistently generate feasible solutions.
Scalability.
Figure 5(d) shows policies trained on $|P|=4$ achieve 80–90% of maximum revenue when deployed on voyages with $|P|\in{5,6}$, demonstrating effective generalization to longer voyages. Inference costs scale linearly with problem sizes, while each solution remains feasible. Note that scaling to larger vessels currently requires retraining.
Deployment Challenges.
Due to limited data availability, we simulate instances matching demand to capacity, as in industry. For industrial deployment, generating representative data and extending differentiable projection to non-convex constraints are key. Finally, integrating MPP with SPP, uptake, and cargo flow management is possible, but it adds significant complexity and decision interdependencies.
Conclusion and Future Directions
We introduce a stochastic dynamic MDP with state-dependent constraints for MPP under demand uncertainty, addressing a class of challenges often overlooked. We propose a general AM policy with unbiased feasibility projections, trained via actor-critic DRL. Our policy uniquely surpasses state-of-the-art baselines by efficiently generating feasible, high-profit solutions that scale and adapt to diverse instances of a complex problem. Although demonstrated on the MPP, our architecture is a general approach for constrained sequential decision-making under uncertainty. It highlights the importance of AI decision support in practical planning, impacting supply chain reliability and sustainability. Future work on non-convex projections, scalability, and new planning problems will tackle real-world AI challenges.
481216202428PPOSAC1,0001{,}0001,2501{,}2501,5001{,}5001,7501{,}750Number of Scenarios, and DRL ModelProfit (\$)SMIP-NASMIP-PI PPO/UVP/CPSAC/UVP/CP
(a) Profit across scenario sizes on test instances.
@@
(b) Computational time across scenario sizes on test instances.
0.10.30.50.70.91,4501{,}4501,5001{,}5001,5501{,}5501,6001{,}600Coefficient of Variation (CV)Profit (\$)PPO/UVP/CPSAC/UVP/CP
(c) Profit across CV levels on generalization instances.
4561,5002,0002,5003, [^0] Number of PortsProfit (\$)PPO/UVP/CPSAC/UVP/CPMax Profit
(d) Profit across voyage lengths on generalization instances.
Figure 5: Sensitivity analyses.
@@
Human oversight is critical in container shipping decision-support tools for fairness, accountability, and transparency. While models enhance decision-making and provide optimized, feasible plans, human judgment remains essential, especially in safety-critical situations like vessel stability and dangerous cargo. Optimizing solely for revenue might inadvertently bias cargo prioritization, impacting customers and voyages. Safeguards are encouraged to mitigate automation bias, allowing users to challenge outputs and ensuring continuous validation and auditing for trust and fairness in cargo placement. Efficient vessel utilization also contributes to reducing fuel consumption and greenhouse gas emissions.
@@ Appendix A MDP of Master Planning Problem
Sets and Parameters
Provided the sets and parameters in Section Domain Preliminaries, we introduce additional subsets of the transport set TR given port $p\in{P}$:
@@
  • Arrival transports: $\textit{TR}^{\textit{ROB}}_{p}=\{(i,j)\!\in\!{P}^{2}\!\mid\!i<p,j>p\}$
@@
For each combination $(i,j,k)$, we associate the expected demand $\mu^{(i,j,k)}$, standard deviation $\sigma^{(i,j,k)}$, TEU per container $\textit{teu}^{(i,j,k)}$, container weight $w^{(i,j,k)}$, and revenue per container $\textit{rev}^{(i,j,k)}$. These are all combined in the instance parameter $\zeta^{(i,j,k)}\;\forall(i,j)\in\textit{TR},k\in K$
@@
Both teu and $w$ are broadcasted to shape $n_{q}=|K|\times|\textit{TR}|$ in the MDP formulations for consistency in dimensionality.
@@
The expected value $\mu\in\mathbb{R}^{n_{q}}_{\geq 0}$ and standard deviation $\sigma\in\mathbb{R}^{n_{q}}_{>0}$ of demand are randomly generated, as will be described in Appendix C.
MDP
We define the MDP by decomposing the traditional CO problem outlined in Van Twiller et al. (2024a).
@@
Algorithm 2 Episode of Decomposed MDP
[…]
@@
$T_{\text{leave port}}$ $=\Big{\{}t\in H\ \mid\exists p\in P^{N_{P}-1}_{1}\text{ such that }$
@@
$T_{\text{new port}}$ $=\Big{\{}t\in H\ \mid\exists p\in P^{N_{P}-1}_{1}\text{ such that }$
@@
Substituting Load Operations for Actions
@@
Feasible Region for Actions
@@
$\textit{teu}^{(\textit{pol}_{t},\textit{pod}_{t},k_{t})}\mathbf{1}^{\top}x_{t}\leq c-u^{\prime}_{t}\textit{teu}\;,$
Stability Constraints
@@
[…]
@@
$\textit{ho}(s,p)$ $=\textit{hm}(s,p)\Bigg{(}\sum_{k\in K}\sum_{\textit{tr}\in\textit{TR}^{\textit{ROB}}_{p}}u_{t}^{(b,d^{\textit{above}},k,\textit{tr})}\Bigg{)}$ (15)
@@
${\textit{cm}}(s,p)$ $=\max\Bigg{(}\sum_{d\in D}\sum_{k\in K}\sum_{\textit{tr}\in\textit{TR}^{M}_{p}}u_{t}^{(0:|B|-1,d,k,\textit{tr})}$
$\qquad\qquad+u_{t}^{(1:|B|,d,k,\textit{tr})}-\overline{\textit{cm}}(s,p),0\Bigg{)}$ (17)
@@ Appendix B Feasibility Mechanisms
Log Probability Adjustments
This subsection provides technical details on the adjustments to the log-probability distribution of the policy as a result of non-linear transformations to distribution samples.
Projecting actions alters the policy’s probability density, necessitating consideration of the change of variables principle (Bishop 2006). This principle ensures valid volume scaling by requiring the transformation $f(x)$ to satisfy:
@@
Weighted Scaling Projection Layer
@@
Unbiased Violation Projection
@@
Finally, we can confirm that Jacobian adjustment is allowed for the unbiased violation projection layer by:
@@
Policy Clipping
We can implement a clipped Gaussian distribution that enforces element-wise bounds on a standard Gaussian (Fujita and Maeda 2018). Let $\mu_{\theta}$ and $\sigma^{2}_{\theta}$ denote the policy’s mean and variance, with bounds $\textit{lb}_{\textit{pc}}$ and $\textit{ub}_{\textit{pc}}$, and $\Phi(\cdot)$ being the cumulative distribution function of the standard Gaussian. Actions are sampled from $\mathcal{N}(\mu_{\theta},\sigma^{2}_{\theta})$ and clipping the result to $\textit{lb}_{\textit{pc}},\textit{ub}_{\textit{pc}}]$. Provided this transformation, we compute the log probabilities $\log\pi(x|s)$ for action $x$ by:
@@
Unbiased Violation Projection
@@
  • $\eta_{v}\in(0,1)$ is a sufficiently small step size parameter to ensure stable convergence.
@@
Since $\mathcal{V}(x)$ is an elementwise projection onto nonnegative values, and some function $h(y)=\max(0,y)$ is convex and non-decreasing, then the function $g(x)=\|\mathcal{V}(x)\|_{2}^{2}$ is convex when AxbAx-b is affine.
@@
By the standard descent lemma (Boyd and Vandenberghe 2004), for a sufficiently small step size $\eta_{v}$, we have:
@@
Since $g(x_{k})$ is also lower-bounded by 0, it must converge to some limit $g^{*}\geq 0$. This implies:
@@ Appendix C Instance Generator
To ensure similar mean and variance of the instances, we derive parameters lb and ub from the definition of the continuous uniform distribution, as follows:
@@ Appendix D Multi-stage Stochastic MIP
Multi-stage Scenario Tree
@@
    A probability measure $P_{\textit{ST}}:V_{\textit{ST}}\to[0,1]$ assigning probabilities to nodes, ensuring:
@@
Intuitively, the scenario tree discretely represents how uncertainty unfolds over time. Each branching point corresponds to a stochastic event, capturing all possible realizations of uncertain parameters across the planning horizon. As we progress through stages, new uncertainty is realized, and decisions are made at each node based on available information. These decisions do not anticipate future realizations, but are optimized over the expected outcomes of all future scenarios encoded in the tree.
MIP Formulation
@@
All integer constraints are relaxed linearly in the implementation. Additionally, we use a sufficiently large constant, denoted by big $M$, to impose logical constraints as needed.
Objective. The objective function (18) maximizes the revenue with parameter $\textit{rev}^{(i,j,k)}\in\mathbb{R}_{>0}$ and minimizes hatch-overstowage with parameter $\textit{ct}^{\textit{ho}}\in\mathbb{R}_{>0}$ and crane moves costs with parameter $\textit{ct}^{\textit{cm}}\in\mathbb{R}_{>0}$ over scenario paths $\phi\in\mathcal{Z}$ each with probability $\mathbb{P}_{\phi}$. We assume each scenario path has equal probability.
@@
max $\sum_{\phi\in\mathcal{Z}}\mathbb{P}_{\phi}\sum_{p\in{P}}\sum_{b\in{B}}\sum_{d\in{D}}\sum_{k\in{K}}\sum_{\textit{tr}\in\textit{TR}^{\textit{+}}(p)}\textit{rev}^{(i,j,k)}\tilde{u}^{b,d,\phi}_{\textit{tr},k}$
$\qquad\qquad\qquad\qquad\qquad-\textit{ct}^{\textit{ho}}\tilde{\textit{ho}}^{\phi}_{p,b}-\textit{ct}^{\textit{cm}}\tilde{\textit{cm}}^{\phi}_{p}$ (18)
@@
$\sum_{k\in{K}}\sum_{\textit{tr}\in\textit{TR}^{\textit{ROB}}(p)}\tilde{u}^{b,d_{o},\phi}_{\textit{tr},k}-M(1-\tilde{\textit{hm}}^{\phi}_{p,b})\leq\tilde{\textit{ho}}^{\phi}_{p,b}$
@@
$\sum_{b\in{b^{\prime}}}\sum_{d\in{D}}\sum_{k\in{K}}\sum_{\textit{tr}\in\textit{TR}^{M}(p)}\tilde{u}^{b,d,\phi}_{\textit{tr},k}-{\overline{z}}^{\phi}_{p}\leq\tilde{\textit{cm}}^{\phi}_{p}$
@@ Appendix E Deep RL Implementation Details
Feasibility Regularization and Primal-Dual RL
To encourage the policy to produce feasible actions with respect to linear constraints $A(s_{t})x\leq b(s_{t})$, we introduce a soft penalty term into the actor loss:
$\mathcal{L}(\theta)$ $=-\mathcal{L}_{\text{actor}}(\theta)+\lambda_{f}\mathcal{L}_{\text{feas}}(\theta),$ (31)
$\mathcal{L}_{\text{feas}}(\theta)$ $=\mathbb{E}_{t}\left[\left(A(s_{t})x_{\theta}(s_{t})-b(s_{t})\right)_{>0}\right],$ (32)
where $(\cdot)_{>0}=\max(0,\cdot)$ denotes the hinge function penalizing constraint violations. The coefficient $\lambda_{f}$ balances reward maximization and constraint satisfaction.
We consider two methods to handle $\lambda_{f}$:
  • Static FR fixes $\lambda_{f}$ via hyperparameter tuning before training.
  • Primal-Dual FR treats $\lambda_{f}$ as a dual variable and jointly learns $\lambda_{f}$ during training by adding a dual output head to the critic network, updated via gradient ascent to enforce constraints adaptively.
This primal-dual approach alleviates the need for manual tuning and improves constraint satisfaction during training, as demonstrated in Ding et al. (2020).
PPO Algorithm
PPO is an on-policy reinforcement learning algorithm that seeks to maximize expected cumulative reward while enforcing stable policy updates via clipped importance sampling (Schulman et al. 2017), as outlined in Algorithm 3. The agent collects trajectories, computing $n_{\text{ppo}}$-step return to evaluate performance with $V_{\theta}(s)$ as estimated state value:
@@
To reduce variance, we adopt Generalized Advantage Estimation (GAE) (Schulman et al. 2016):
@@
SAC Algorithm
Soft Actor-Critic (SAC) is an off-policy reinforcement learning algorithm that optimizes both reward maximization and entropy to encourage efficient exploration (Haarnoja et al. 2018), as outlined in Algorithm 4. It is based on maximum entropy reinforcement learning, which aims to learn a stochastic policy that not only maximizes cumulative rewards but also maintains high entropy for robustness and stability. SAC leverages a soft Q-learning approach, using two Q-functions to mitigate overestimation bias, an entropy-regularized policy update, and an automatically adjusted temperature parameter to balance exploration and exploitation.
@@
1:for each iteration do
2:  for each environment step $t$ do
3:   Sample action $x_{t}\sim\pi_{\theta}(x_{t}|s_{t})$
4:   Perform transition $s_{t+1}\sim\mathcal{T}(s_{t+1}|s_{t},x_{t})$
5:   Observe reward $r_{t}=\mathcal{R}(s_{t},x_{t})$,
6:   Store $(s_{t},x_{t},r_{t},s_{t+1})$ in $\mathcal{D}$.
7:  end for
8:  for each gradient step do
9:   Sample a minibatch $(s_{t},x_{t},r_{t},s_{t+1})$ from $\mathcal{D}$.
10:   Compute target Q-value: $Q_{\text{target}}(s_{t},x_{t})$
11:   Update parameters:  $\theta_{\text{critic}}^{l}\leftarrow\theta_{\text{critic}}^{l}-\eta_{c}\nabla_{l}\mathcal{L}_{\text{critic}}(\theta)\text{ for }l\in\{1,2\}$$\theta_{\text{actor}}\leftarrow\theta_{\text{actor}}-\eta_{a}\nabla\mathcal{L}_{\text{actor}}(\theta)$$\alpha\leftarrow\alpha-\eta_{\alpha}\nabla\mathcal{L}_{\alpha}(\theta)$$\theta_{\text{target}}^{l}\leftarrow\tau\theta_{\text{critic}}^{l}+(1-\tau)\theta_{\text{target}}^{l}\text{ for }l\in\{1,2\}$
12:  end for
13:end for
14:return  Policy $\pi_{\theta}$
Hyperparameters
The parameters of the MPP instances, presented in Table 3, are chosen based on domain knowledge to generate problem instances that closely reflect real-world scenarios.
@@
Table 3: Instance parameters.
Table 4 provides the hyperparameters of projected and vanilla PPO and SAC. Hyperparameter tuning is performed using Bayesian optimization via the Weights & Biases platform (Biewald 2020). Only a selected subset of hyperparameters are optimized, each bounded by specified minimum and maximum values in Table 4. The tuning process uses the objective value as the primary selection criterion, prioritizing feasible solutions whenever possible.
Settings Projection Algorithms Vanilla Algorithms Ranges
Hyperparameters Symbol PPO SAC PPO SAC Min Max
@@
Number of Heads 8 8 4 4 4 16
Hidden Layer Size 128 128 256 256 128 512
Encoder Layers 3 3 2 1 1 3
Decoder Layers 3 3 3 3 1 3
@@
Dropout Rate $0.009$ $0.009$ $0.073$ $0.164$ $0.001$ $0.2$
Max Policy Std. $1.931$ $1.931$ $1.118$ $1.779$ $0.01$ $10.0$
@@
Learning Rate (LR) $\eta$ $1.47\times 10^{-4}$ $1.47\times 10^{-4}$ $9.64\times 10^{-4}$ $9.10\times 10^{-4}$ $1.0\times 10^{-6}$ $0.01$
Primal-Dual LR $\eta_{\mathit{pd}}$ - - $3.47\times 10^{-5}$ $3.47\times 10^{-5}$ $1.0\times 10^{-6}$ $0.01$
Batch Size 64 64 64 64 16 512
Embedding Size 128 128 128 128 64 256
@@
Feasibility Penalty $\lambda_{f}$ Vector Vector Vector Vector $0.01$ $10.0$
@@
Mini-batch Size 32 32 32 32 16 64
Update Epochs 5 1 5 1 3 20
Entropy Target N/A |X|-|{X}| N/A |X|-|{X}|
Projection LR $\eta_{v}$ $0.01$ $0.01$ N/A N/A $0.001$ $0.1$
Projection Epochs $300$ $300$ N/A N/A 50 500
Inference Threshold $\delta_{v}$ $0.01$ $0.01$ N/A N/A $0.001$ $0.1$